prettygirlLA:

The area of a triangular sail is given by the expression 1 2 bh, where b is the length of the base and h is the height. What is the area of a triangular sail in a model sailboat when b = 14 inches and h = 8 inches?

Serenity1jackson:

um

Serenity1jackson:

just go to brainly

prettygirlLA:

k

Serenity1jackson:

umhm no prob

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Serenity1jackson just go to brainly \(\color{#0cbb34}{\text{End of Quote}}\) We would prefer that you don't encourage users to go to another help site when we can surely help them out here.

XioGonz:

@prettygirlLA Do you have a picture?

prettygirlLA:

no picture was shown

prettygirlLA:

but the answer is suppose to ___in to the power of 2

b3lla2006:

10 I think

jhonyy9:

so with this formula i think you wan say \[A = \frac{ 1 }{ 2 }bh\]

b3lla2006:

1/2 x 12 x 7= 42

jhonyy9:

The area of a triangular sail is given by the expression 1 2 bh, where b is the length of the base and h is the height. What is the area of a triangular sail in a model sailboat when b = 14 inches and h = 8 inches? so i dont see there 12 and 7 ?

b3lla2006:

the area is 42

jhonyy9:

b = 14 inches h = 8 inches \[A = \frac{ 1 }{ 2 }bh\]

b3lla2006:

sorry 56

prettygirlLA:

yup you were right 56in to the powerof 2

prettygirlLA:

thank for the help

jhonyy9:

np yw

b3lla2006:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @prettygirlLA thank for the help \(\color{#0cbb34}{\text{End of Quote}}\) np

jhonyy9:

@XioGonz

XioGonz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 b = 14 inches h = 8 inches \[A = \frac{ 1 }{ 2 }bh\] \(\color{#0cbb34}{\text{End of Quote}}\) I agree!